### v1

# An operational approach to classifying measurement incompatibility

Abstract
Measurement incompatibility has proved to be an important resource for information processing tasks. In this work, we analyze various levels of incompatibility of measurement sets. We provide operational classification of measurement incompatibility with respect to two elementary classical operations, viz., coarse-graining of measurement outcomes and convex mixing of different measurements. We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or convex mixing. Robustness against white noise is investigated for mutually unbiased bases that can sustain full incompatibility. Furthermore, we propose operational witnesses for different levels of incompatibility subject to classical operations, using the input-output statistics of Bell-type experiments as well as experiments in the prepare-and-measure scenario.
I introduction
Measurement incompatibility is a concept relating observables which cannot be measured jointly with arbitrary accuracy Gühne et al. (2023). It is a purely quantum effect of which the most well-known example concerns the position and momentum of a quantum particle. Being a fundamental concept of quantum theory, it takes a pivotal role in explaining several quantum phenomena, such as, Bell-nonlocality Fine (1982a); Wolf et al. (2009), Einstein–Podolsky–Rosen steering Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015), measurement uncertainty relations Pramanik et al. (2013); Busch et al. (2014); Maity et al. (2017); Saha et al. (2020), quantum contextuality Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), quantum violation of macrorealism Mal and Majumdar (2016); Uola et al. (2019), and temporal and channel steering Karthik et al. (2015); Banik et al. (2015); Uola et al. (2018). Measurement incompatibility finds applications in various quantum information processing tasks. Recently, it has been shown that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al. (2023).
The significance of measurement incompatibility in various quantum effects calls for its in-depth characterization. Towards this direction, a classification of measurement incompatibility with respect to projection onto subspaces has been recently performed Uola et al. (2021). In the present work, our aim is to classify measurement incompatibility with respect to certain basic classical operations. From an operational perspective, here we employ two such operations, namely coarse-graining of measurement outcomes and convex mixing of measurement settings. Coarse-graining of measurement outcomes arises naturally Toscano et al. (2018) in several instances, such as, for example, quite obviously in measurements on continuous variable system. Though the eigen spectra of the observables are infinite-dimensional taking continuous values, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values. This inaccuracy in the recording of measurement outcomes is manifested in the coarse-graining of measurement outcomes, which is inevitable in practice. Coarse-graining has also been employed to study the phenomenon of quantum to classical transition. It is observed that quantum phenomena may disappear due to imprecision of measurement outcomes Kofler and Brukner (2007); Rudnicki et al. (2012); Jeong et al. (2014); Mal and Majumdar (2016).
Device imperfection may also lead to the measurement device performing probabilistically a set of measurements, instead of always performing the desired particular measurement. In such a case, a convex mixing of the given set of measurements arises effectively. As incompatibility of measurements is a quantum concept, it is interesting to examine how this property behaves under such elementary classical operations. Our study is motivated by the question as to how one may compare the degree of incompatibility between two different sets of measurements. For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one. A similar argument holds for the case of convex-mixing of measurements too.
As incompatible measurements are useful for performing various information processing tasks, given a measurement device claiming to produce incompatible measurements, we must certify it before using it in an information processing task. Verification of incompatiblity of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning. Device-independent protocols are conceptually most powerful, relying only on the input-output statistics Brunner et al. (2014), with a wide range of applications from quantum cryptography Ekert (1991); Acín et al. (2007) to communication complexity Buhrman et al. (2010), self-testing Mayers and Yao (2004); Šupić and Bowles (2020); Goswami et al. (2018); Bian et al. (2020) and incompatibility witnesses Quintino et al. (2019). However, they require shared entanglement, an expensive resource, and prohibition of communication between the involved parties. On the other hand, semi-device-independent protocols are inspired by the standard prepare-and-measure scenario, with an additional constraint on the dimension of the communicated quantum states Van Himbeeck et al. (2017); Moreno et al. (2021). They have been applied in quantum key distribution Pawłowski and Brunner (2011), quantum networks Bowles et al. (2015), randomness certification Passaro et al. (2015), random access codes Li et al. (2012), and certification of meassurements Maity et al. (2021); Das et al. (2022).
In the present work we investigate the issue of certification of measurement incompatibility subjected to the classical operations of coarse-graining and convex mixing of measurements from both the device-independent and the semi-device-independent perspective. An essential ingredient in the performance of quantum information protocols is the presence of noise. Within the context of our present study, noise is reflected in the degrading of incompatibility properties of various measurement sets. We examine the critical amount of robustness of the noisy measurement above which the measurements remain incompatible under the mentioned classical operations. We then study incompatibility witnesses in the device-independent and the semi-device-independent scenarios.
Our paper is structured as follows. In Sec. (II) we present the basic mathematical ingredients for classifying measurement incompatibility under the two operations of coarse-graining of outcomes and convex-mixing of measurements, respectively. In Sec. (III) we study the robustness under noise for various levels of incompatibility subjected to the above classical operations. In Sec. (IV) we define operational witnesses of incompatibility, and furnish examples to study their performance in both device-independent and the semi-device-independent frameworks. Concluding remarks are presented in Sec. (V).
II Classifying measurement incompatibility under elementary classical operations
In the case of projective measurements, the observables whose corresponding operators commute are jointly measurable. However, this is not true in the case of general quantum measurements. The most general quantum measurement is described by Positive Operator Valued Measures (POVM), which is a set of operators, $\{O_{\alpha}\}$ with $0\leq O_{\alpha}\leq\mathbbm{1}$, and $\sum_{\alpha}O_{\alpha}=\mathbbm{1}$. A set of measurements defined by $\{M_{z_{x}|x}\}$ (here $x$ corresponds to the measurement, $z_{x}$ denotes the outcomes of the measurement labeled by $x$) is compatible if there exists a parent POVM, $G_{\lambda}$ and classical post-processing for each $x$ given by $\{p(z_{x}|x,\lambda)\}$ such that
$\forall z_{x},x,\quad M_{z_{x}|x}=\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda},$ (1)
where $0\leq p(z_{x}|x,\lambda)\leq 1,\sum_{z_{x}}p(z_{x}|x,\lambda)=1\forall x,\lambda$    Gühne et al. (2023).
II.1 Coarse-graining of outcomes
Coarse-graining of the outcomes of a measurement is a concept where some of the outcomes are treated on the same footing, i.e., different outcomes are clubbed together to form a single outcome. After the coarse-graining of the outcomes, the effective number of outcomes reduces. Coarse-graining may occur in practice due to measurement errors. Additionally, coarse-graining is used in when one does not require a finer description of the values of the measurement outcomes, but rather a broad description suffices to understand the nature of some physical property.
Let us introduce the notation $[k]:=\{0,\cdots,k-1\}$ for any natural number $k.$ Consider a $\overline{d}$ outcome measurement $\{M_{\overline{z}}\}$ and an arbitrary coarse-graining that yields a $d$ outcome measurement $\{M_{z}\}$, where $\overline{z}\in[\overline{d}]\,\,\text{and}\,\,z\in[d]$. The coarse-graining can be defined by conditional probabilities $\{c(z|\overline{z})\}$ such that
$M_{z}=\sum_{\overline{z}}c(z|\overline{z})M_{\overline{z}},$ (2)
and $c(z|\overline{z})\in\{0,1\}$. Trivial coarse-graining is the one where the measurement becomes single outcome, that is, $c(z|\overline{z})=1,\forall\overline{z}$. Two incompatible measurements may not necessarily remain incompatible after certain coarse-graining of their outcomes.
Definition 1 (Fully incompatible measurements w.r.t coarse-graining).
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is fully incompatible with respect to (w.r.t) coarse-graining if they remain incompatible after all possible nontrivial coarse-graining. That is, if the resultant set of measurements given by
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},$ (3)
after all possible sets of nontrivial coarse-graining $\{c_{x}(z_{x}|\overline{z}_{x})\}$ pertaining to every setting $x$ is incompatible, then we call them fully incompatible.
Definition 2 ($d$-incompatible measurements w.r.t coarse-graining).
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is $d$-incompatible w.r.t coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that give rise to at least $d$ outcome measurements. In other words, if the resultant set of measurements
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},\ z_{x}\in[k_{x}],k_{x}\geqslant d,\forall x$ (4)
after all possible coarse-graining resulting in at least $d$ outcome measurements is incompatible, then we call them $d$-incompatible.
For example, the four outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
$M=\Big{\{}|0\rangle,|1\rangle,|2\rangle,|3\rangle\Big{\}}$ (5)
and
$N=\Big{\{}|0+1\rangle,|0-1\rangle,|2+3\rangle,|2-3\rangle\Big{\}}$ (6)
is 3-incompatible, where the notation $|0+1\rangle=|0\rangle+|1\rangle,$ and similarly for others.
Observation 1.
A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t coarse-graining.
If a set of measurements is 2-incompatible, then it implies that the set remains incompatible after all possible coarse-graining of the outcomes such that the number of outcomes of each measurement in the newly formed set of measurements is greater than or equal to two. Also, the lowest number of outcomes of measurement is two for a nontrivial coarse-graining. Furthermore, if a set of measurements is $d$-incompatible, then, by definition, it is $n$-incompatible as well, where $n>d$, but the reverse is not true. This proves Observation 1.
Observation 2.
[…]
$\left[\sum_{i\in\mathcal{M}_{k}}P_{i},\sum_{j\in\mathcal{N}_{l}}Q_{j}\right]\neq 0,\ \forall k,l.$ (7)
Proof.
[…]
Theorem 1.
The following condition is necessary but not sufficient for two rank-one projective measurements, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t coarse-graining:
$\langle\psi_{i}|\phi_{j}\rangle\neq 0,\ \forall i,j.$ (8)
The above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
Proof.
First, note that for sharp measurement compatibility and commutativity are equivalent Heinosaari and Wolf (2010). Suppose $\exists i,j,$ such that $\langle\psi_{i}|\phi_{j}\rangle=0$. Consider coarse-graining of all other outcomes except $i$ and $j$ for the two measurements. Then, the resulting measurements will be $\{|\psi_{i}\rangle\!\langle\psi_{i}|,\mathbbm{1}-|\psi_{i}\rangle\!\langle\psi_{i}|\}$ and $\{|\phi_{j}\rangle\!\langle\phi_{j}|,\mathbbm{1}-|\phi_{j}\rangle\!\langle\phi_{j}|\}$, which commute with each other (i.e., the resulting measurements are compatible) since $\langle\psi_{i}|\phi_{j}\rangle=0$. Thus, if the measurements are fully incompatible, condition (8) holds.
To show that (8) is not sufficient, consider the following two $4$-dimensional rank-one projective measurements (with the normalization factor $1/\sqrt{2}$),
$\{|0\rangle+|1\rangle,|0\rangle-|1\rangle,|2\rangle+|3\rangle,|2\rangle-|3\rangle\}$
$\{|0\rangle+|2\rangle,|0\rangle-|2\rangle,|1\rangle+|3\rangle,|1\rangle-|3\rangle\}.$ (9)
One can check that (8) holds for all $i,j=0,1,2,3$. But coarse-graining of outcomes $0,1$ and $2,3$ for both the measurements leads to
$\{|0\rangle\!\langle 0|+|1\rangle\!\langle 1|,|2\rangle\!\langle 2|+|3\rangle\!\langle 3|\},$
$\{|0\rangle\!\langle 0|+|2\rangle\!\langle 2|,|1\rangle\!\langle 1|+|3\rangle\!\langle 3|\},$ (10)
which are compatible.
[…]
II.2 Convex mixing of measurements
The concept of convex mixing of measurements may be best understood by considering an example: suppose we have a measurement device that produces three different measurements $(M_{0},M_{1}\,\,\text{and}\,\,M_{2})$ and has two input switches. If the first switch is ON it performs measurement $M_{0}$ with probability $q$ and $M_{1}$ with probability $(1-q)$ and for the second switch, it always performs $M_{2}.$ Thus, in principle, the device produces two measurements: $M_{2}$ and a convex mixture of $M_{0}$ and $M_{1}.$
Suppose $M_{0}=\{M_{z|0}\},M_{1}=\{M_{z|1}\},M_{2}=\{M_{z|2}\}$ such that they have same number of outcomes $z\in[d]$. Now the measurement that is realized by a convex combination of $M_{0}$ and $M_{1}$ with some weightage $q\,\,\text{and}\,\,1-q$, where $q\in[0,1]$is defined as
$M_{(0,1,\pi)}=\{qM_{z|0}+(1-q)M_{\pi(z)|1}\},$ (11)
where $\pi(z)$ is any permutation on outcome $z$. Here $\pi(z)$ is a bijective function from the set $z\in[d]$ to itself. Even if $M_{2}$ is incompatible with $M_{0}$ and $M_{1}$ separately, $M_{2}$ is not necessarily incompatible with $M_{(0,1,\pi)}$ for all values of $q$.
Definition 3.
[…]
Consider three unbiased qubit measurements,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\vec{n_{i}}\cdot\vec{\sigma}\right),\quad i=0,1,2,$ (12)
and $||\vec{n_{i}}||\leqslant 1$. These three measurements are fully incompatible w.r.t convex mixtures if and only if (see eq.(7) of Gühne et al. (2023)),
$||\vec{n}_{i}+q\vec{n}_{j}\pm(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}\mp(1-q)\vec{n}_{k}||>2,$ (13)
for all $q$, and for all $(i,j,k)\in\{(0,1,2),(1,2,0),(2,0,1)\}$.
Observation 3.
If three qubit measurements (12) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t convex mixtures.
Proof.
Note that permutation of the outcomes of $M_{i}$ in (12) is nothing but taking $-\vec{n}_{i}$ instead of $\vec{n}_{i}$. If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ and permutations of outcomes such that
$\vec{n}_{0}=\frac{1}{c}(q\vec{n}_{1}+(1-q)\vec{n}_{2})$ (14)
for some $c\in(0,1],q\in[0,1]$. In other words, there exists a triple $(i,j,k)$ and permutation of outcomes such that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), such that sum of those two non-negative coefficients is greater than or equal to 1. Substituting this into left hand side of (13) taking $+$ sign, we find
$||(1+c)\vec{n}_{i}||+||(1-c)\vec{n}_{i}||\leqslant|1+c|+|1-c|=2.$ (15)
[…]
Observation 4.
Suppose three qubit measurements (12) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leq\nu_{1},\nu_{2},\nu_{3}\leq 1$, that is, they are the noisy version of Pauli observables,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\nu_{i}\sigma_{i}\right)=\nu_{i}\left(\frac{\mathbbm{1}\pm\sigma_{i}}{2}\right)+(1-\nu_{i})\frac{\mathbbm{1}}{2},$ (16)
taking $i=0,1,2$ and $(\sigma_{0},\sigma_{1},\sigma_{2})=(\sigma_{x},\sigma_{y},\sigma_{z})$. Then they are fully incompatible w.r.t convex mixture if and only if
$\min\bigg{\{}\nu^{2}_{0}+\frac{\nu^{2}_{1}\nu^{2}_{2}}{\nu^{2}_{1}+\nu^{2}_{2}},\nu^{2}_{1}+\frac{\nu^{2}_{0}\nu^{2}_{2}}{\nu^{2}_{0}+\nu^{2}_{2}},\nu^{2}_{2}+\frac{\nu^{2}_{0}\nu^{2}_{1}}{\nu^{2}_{0}+\nu^{2}_{1}}\bigg{\}}>1.$ (17)
Proof.
In terms of $\nu_{i}$, the left hand side of (13) becomes $2\sqrt{\nu^{2}_{i}+q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}}$. Note that the minimum of $q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}$ occurs at $\tilde{q}=\nu_{k}^{2}/(\nu^{2}_{j}+\nu^{2}_{k})$. Since $0\leq\nu_{k}^{2}/(\nu^{2}_{j}+\nu^{2}_{k})\leq 1$ for any $\nu_{j},\nu_{k}\in[0,1]$, the above-mentioned minimum can always be achieved with a suitable choice of $q$. Hence, we have that
$2\sqrt{\nu^{2}_{i}+q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}}$ $\geq 2\sqrt{\nu^{2}_{i}+\tilde{q}^{2}\nu^{2}_{j}+(1-\tilde{q})^{2}\nu^{2}_{k}}$
$=2\sqrt{\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}}$ (18)
Thus, the r.h.s of (13) is greater than 2 for all values of $q$ if and only if
$\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}>1.$ (19)
[…]
Clearly, the three Pauli observables are fully incompatible w.r.t. convex mixture, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (17) implies $\nu>\sqrt{2/3}$.
We can generalize this notion for a set of $n$ measurements.
Definition 4 ($k$-incompatible measurements w.r.t. convex mixtures).
A set of measurements is $k$-incompatible w.r.t convex mixtures if after taking all possible convex mixtures over all possible $k$ partitions the resulting $k$ number of measurements from the set of $n$ measurements (with relabelling of outcomes) is incompatible.
Definition 5 (Fully incompatible measurements w.r.t. convex mixtures).
A set of measurements is fully incompatible w.r.t convex mixtures if it remains incompatible after taking all possible convex mixtures over all possible partitions of $n$ measurements (with relabelling of outcomes).
Theorem 2.
Fully incompatible measurements w.r.t convex mixtures imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
Consider a set of $n$ measurements, $M=\{M_{1},M_{2},M_{3},\ldots,M_{n}\}$, in which there is a pair of measurements $M_{i}$ and $M_{j}$ which are compatible. Now if we make two partitions where $M_{i}$ and $M_{j}$ are in different partitions, the convex mixture of measurements must be compatible. This is true for any compatible pair $M_{i}$ and $M_{j}$. Thus, if the measurements are fully incompatible w.r.t convex mixture, every pair of measurements must necessarily be incompatible.
The reverse is not true. Consider the three noisy Pauli measurements of Eq.(16). It can be shown by using semi-definite programming that if $0.71<\nu\leq 0.81$, the measurements are pairwise incompatible but it is not fully incompatible w.r.t convex mixture.
III Robustness under noise for different levels of incompatibility
In this section, we analyze the role of noise on quantum measurements and study how the incompatibility properties depend on it. Due to the ubiquitous nature of noise, it is pertinent to study the extent to which noise could be tolerated by a set of measurements while still retaining their incompatibility. We take a noisy version of mutually unbiased bases measurements, $M_{i|x}\equiv\{\nu|\phi_{i}^{x}\rangle\langle\phi_{i}^{x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d}\},\,\,\text{where}\,\,\{|\phi_{i}^{x}\rangle\}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$. Here $\nu$ is the robustness parameter (or visibility parameter) and $(1-\nu)$ is the noise parameter, $0\leq\nu\leq 1$. When the noise parameter is zero (i.e., robustness, $\nu=1$), the measurements are fully incompatible, and when the noise parameter is one (i.e., the robustness, $\nu=0$), the measurements are trivial and compatible. Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible.
To check the compatibility, i.e., the existence of a parent POVM, we use the method described in Gühne et al. (2023). This can be casted as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
$\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda}=M_{z_{x}|x}\forall x,z_{x},$ (20)
$\sum_{\lambda}G_{\lambda}=\mathbbm{1},$ (21)
$G_{\lambda}\geq\mu\mathbbm{1},$ (22)
where $\mu$ is the optimization parameter. This method finds the maximum value of $\mu$ for each $\{p(z_{x}|x,\lambda)\}$. If this optimization returns a negative value of $\mu$, then the constraint of Eq.(22) can’t be fulfilled, which implies that the measurements $\{M_{z_{x}|x,\lambda}\}$ are incompatible. Otherwise, they are compatible.
III.1 Coarse-graining of outcomes
Dimension $3$. Let $\{P_{i}\}$ and $\{Q_{j}\},i,j\in[3]$ be two measurements acting on $\mathbb{C}^{3}$, where
$P_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
$Q_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (23)
with
$|\psi_{0}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+|1\rangle+|2\rangle),$
$|\psi_{1}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega|1\rangle+\omega^{2}|2\rangle),$
$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$
It is observed that for $\nu>0.683$, the two measurements are incompatible. For finding ”$2$-incompatible w.r.t coarse-graining”, we consider all possible coarse-graining of outcomes of the two measurements such that after coarse-graining, the new set of measurements have at least two outcomes each. We find that these measurements are incompatible w.r.t all possible coarse-graining for $\nu>0.711$. Note that for $0.683<\nu\leq 0.711$, the incompatible measurements become compatible under certain coarse-graining of outcomes (see TABLE 1).
Dimension $4$. For checking incompatibility w.rt coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated for two POVM measurements with $4$ outcomes each. The corresponding measurements are $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\},i,j\in[4]$.
$P^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
$Q^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (24)
with
$|\psi^{\prime}_{0}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle+|2\rangle+|3\rangle),$
$|\psi^{\prime}_{1}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle-|2\rangle-|3\rangle),$
$|\psi^{\prime}_{2}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle-|2\rangle+|3\rangle),$
$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle),$ (25)
for $\nu>0.666$, these measurements are incompatible. We find that for $\nu>0.675$, the measurements are incompatible under all possible non-trivial coarse-graining that gives rise to at least $3$ outcomes. The measurements are fully incompatible for $\nu>0.72$. The results are summarized in TABLE 1.
In dimension $3$ In dimension $4$
Incompatible 0.683 0.666
2-incompatible 0.711 0.720
3-incompatible 0.683 0.675
Table 1: Critical values of robustness for MUBs w.r.t coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
III.2 Convex mixing of measurements
Dimension $3$. For checking incompatibility w.r.t convex mixtures (CM), a minimum of three measurements are needed. The two measurements are $\{P_{i}\}$ and $\{Q_{j}\}$, as given in Eqs.(23), The third measurement is $\{R_{k}\},k\in[3]$ where
$R_{k}=\nu|\phi_{k}\rangle\langle\phi_{k}|+(1-\nu)\frac{\mathbbm{1}}{3}.$ (26)
with
$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},|\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}},\end{split}$ (27)
$\omega$ being the cube roots of unity.
The results are summarized in TABLE 2.
Dimension $4$. Consider the measurements $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\}$ as defined in Eqs.(III.1) and(III.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in[4]$.
$\{R^{\prime}_{k}\}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (28)
with
$|\phi^{\prime}_{0}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle-\mathbbm{i}|2\rangle-\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{1}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+\mathbbm{i}|2\rangle+\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{2}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle+\mathbbm{i}|2\rangle-\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle-\mathbbm{i}|2\rangle+\mathbbm{i}|3\rangle).$ (29)
The results are summarized in TABLE 2.
In dimension $3$ In dimension $4$
Incompatible 0.537 0.692
2-incompatible 0.764 0.705
3-incompatible 0.537 0.692
Table 2: Critical values of robustness for MUBs w.r.t convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
IV Operational witnesses of incompatibility
IV.1 Coarse-graining of outcomes
As we mentioned earlier, measurements that are fully incompatible w.r.t coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t coarse-graining. To operationally certify whether a set of measurements are fully incompatible w.r.t coarse-graining is important from the perspective of determing the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al. (2009); Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015); Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), as well as for checking the proficiency of such a set for information processing tasks Carmeli et al. (2020); Saha et al. (2023). Operational certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning. Below we study the two classes of operational certification separately
Device-independent witness
Here one requires neither any prior knowledge of the internal functioning of the measurement device, nor any idea of the dimension of the system on which the measurements act. In Wolf et al. (2009), it has been proven that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality. We can use this result to witness incompatibility for any two 3-outcome projective measurements before and after coarse-graining, from the input-output statistics.
Semi-Device-independent witness
Here, although we do not have any prior knowledge of the internal functioning of the measurement device, we do know the dimension of the system on which the measurements act. Any two incompatible rank-one projective measurements provide an advantage in $(2,d,d)-$RAC Saha et al. (2020). If any two POVMs each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average maximum success probability in $(2,\overline{d},d)$ RAC is bounded by Saha et al. (2023)
$P(2,\overline{d},d)\leq\frac{1}{2}(1+\frac{d}{\overline{d}^{2}})=P_{CB}(2,\overline{d},d),$ (30)
where $P_{CB}(2,\overline{d},d)$ is the maximum average success probability obtained in a $(2,\overline{d},d)$ RAC using two compatible measurements. For full incompatibility of two 3-outcome rank-one projective measurements, we prove the following:
Theorem 3.
Two rank-one projective measurements, $M=\Big{\{}|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\Big{\}}$ and $N=\Big{\{}|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\Big{\}}$, can be witnessed to be fully incompatible w.r.t. coarse-graining via RAC iff $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
Proof.
Without loss of generality, any pair of three outcome rank-one projective measurements can be written up to unitary freedom as,
$M=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\Big{\}}\equiv\Big{\{}|\phi_{i}\rangle\langle\phi_{i}|\Big{\}},$
and
$N=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}\equiv\Big{\{}|\psi_{j}\rangle\langle\psi_{j}|\Big{\}},$ (31)
where $|\psi_{j}\rangle=\alpha_{j}|0\rangle+\beta_{j}|1\rangle+\gamma_{j}|2\rangle$.
Let us first coarse grain the second and third outcomes. So, the new measurement pair becomes:
$M^{(2,3)}=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|+|2\rangle\langle 2|\Big{\}},$ (32)
$N^{(2,3)}=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}.$
The new compatibility bound in this case is,
[…]
and the bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in the (2,2,3) RAC if:
$\sum_{x,y=0}^{1}$ $||M^{(2,3)}(x)+N^{(2,3)}(y)||>7,\,\,\text{which implies}$
$||A||$ $+$ $||B||+||C||+||D||>7$ (33)
$\text{with}\,\,A$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(0)$
$=$ $|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Now, $|\psi_{1}\rangle$ can be written as:
$|\psi_{1}\rangle=\langle 0|\psi_{1}\rangle|0\rangle+\langle u|\psi_{1}\rangle|u\rangle\,\,\text{with}$ (34)
$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1.$
$A$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
$A=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (37)
So, the maximum eigenvalue of $A\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
$||A||=1+|\langle 0|\psi_{1}\rangle|.$
By Eq.(IV.1) we have
$B$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(1)$ (38)
$=$ $|0\rangle\langle 0|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|$
$=$ $\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $B$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$B=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (43)
and
$||B||=2+|\langle 0|\psi_{1}\rangle|.$
Also, from Eq.(IV.1)
$C$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(0)$ (44)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{1}\rangle\langle\psi_{1}|\,\,$
$=$ $\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $C$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$C=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (49)
and
$||C||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$
Similarly, from Eq.(IV.1)
$D$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(1)$ (50)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\,\,$
$=$ $2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
and $D$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$D=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix}\,\,\text{where}\,\,\Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (55)
and $||D||=2.$
[…]
Device-independent witness of full-incompatibility of two measurements w.r.t coarse-graining from a single experiment
We now study the following problem: given two measurements of outcomes $d_{1}$ and $d_{2}$ respectively, how can one witness the full-incompatibility w.r.t coarse-graining in a device-independent way from a single experiment? Here ”single experiment” implies the Bell test which gives a Bell-inequality violation for the two measurements of outcomes $d_{1}$ and $d_{2}$ respectively as chosen as inputs of Alice and some measurement choices for Bob and a shared entangled state. As we know if two measurements are $2-$ incompatible w.r.t coarse-graining then they are fully-incompatible w.r.t coarse-graining. Also, we use the result of Wolf et al. (2009), that any two binary outcome incompatible measurements violate the Bell-CHSH inequality. Thus, from the original measurement statistics we do all possible coarse-grainings to make it two outcome measurements and check the Bell-CHSH inequalities. If, all possible coarse-grained two-outcome measurements violate the Bell-CHSH inequalities then the measurements are fully-incompatible w.r.t coarse-graining.
To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the CGLMP inequality Collins et al. (2002), a well-studied Zohren and Gill (2008); Das et al. (2019); Karczewski et al. (2022) Bell-type inequality. The projective measurements of Alice are: $A_{a}\equiv\{|\xi\rangle_{A,a}\},$ where
$|\xi\rangle_{A,a}=\frac{1}{\sqrt{3}}\sum_{j=0}^{2}\text{exp}\bigg{(}\mathbbm{i}\frac{2\pi}{3}j(\xi+\alpha_{a})\bigg{)}|j\rangle_{A},$ (56)
with $a\in\{1,2\}$ corresponding to two different measurement settings of Alice, and $\xi\in\{0,1,2\},\alpha_{1}=0,\alpha_{2}=\frac{1}{2}$. The projective measurements of Bob are: $B_{b}\equiv\{|\eta\rangle_{B,b}\},$ where
$|\eta\rangle_{B,b}=\frac{1}{\sqrt{3}}\sum_{j=0}^{2}\text{exp}\bigg{(}\mathbbm{i}\frac{2\pi}{3}j(-\eta+\beta_{b})\bigg{)}|j\rangle_{B},$ (57)
with, $b\in\{1,2\}$ corresponding to two different measurement settings of Bob, and $\eta\in\{0,1,2\},\beta_{1}=\frac{1}{4},\beta_{2}=-\frac{1}{4}.$
The incompatibility of the measurements of Eq.(56) before coarse-graining is guaranteed by the CGLMP-inequality violation. Now we investigate the incompatibility status of the measurements of Alice after all possible coarse-graining from the measurement statistics obtained in the CGLMP experiment. Alice’s measurements now have two outcomes each and it is a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes. For this scenario, Collins and Gisin have shown that there are a total of 72 CH-facet inequalities Collins and Gisin (2004). The CH-inequalities Fine (1982b) are of the form:
$-1\leq$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (58)
$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2})\leq 0.$
The other CH-inequalities are obtained by (1) interchanging $A_{1}$ with $A_{2}$, (2) interchanging $B_{1}$ with $B_{2}$, and (3) interchanging both $A_{1}$ with $A_{2}$ and $B_{1}$ with $B_{2}$. Consider the scenario where we coarse-grain (0,1) outcomes for both the inputs of Alice. Lets make the following relabeling $(0,1)\equiv\overline{0}$ and $2\equiv\overline{1}$ for the outcomes of $A_{1}$ and $A_{2}$ and also consider the clubbing of $(0,1)\equiv\overline{0}$ outcomes both for $B_{1}$ and $B_{2}.$ Under this relabelling Eq.(58) takes the form:
$-1\leq$ $P(\overline{0}\,\,\overline{0}|A_{1},B_{1})+P(\overline{0}\,\,\overline{0}|A_{1},B_{2})+P(\overline{0}\,\,\overline{0}|A_{2},B_{2})$ (59)
$-P(\overline{0}\,\,\overline{0}|A_{2},B_{1})-P(\overline{0}|A_{1})-P(\overline{0}|B_{2})\leq 0,$
where,
$P(\overline{0}\,\,\overline{0}|A_{i},B_{j})=P(00|A_{i},B_{j})+P(01|A_{i},B_{j})$
$+P(10|A_{i},B_{j})+P(11|A_{i},B_{j}),$ (60)
$P(\overline{0}|A_{i})=P(0|A_{i})+P(1|A_{i}),$
$P(\overline{0}|B_{i})=P(0|B_{i})+P(1|B_{i}),\,\,i,j\in\{1,2\}.$ (61)
Similarly, there are other 8 possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(59). One can check that when there are the same coarse-graining of outcomes for both measurement inputs of Alice, we get a CH-inequality violation. Thus, under these coarse-grainings, the two measurements of Alice remain incompatible. When there are different coarse-grainings, for some cases we get CH-violation, but for other cases, we do not get CH-violation. CH-violation under a particular coarse-graining signifies that the measurements are incompatible. However, when there is no CH-violation we can not conclude anything regarding the incompatibility status. This is depicted in TABLE   3.
$A_{1}$ $A_{2}$ Incompatible
(0,1) (0,1) yes
(0,1) (1,2) ?
(0,1) (0,2) yes
(1,2) (0,1) yes
(1,2) (1,2) yes
(1,2) (0,2) ?
(0,2) (0,1) ?
(0,2) (1,2) yes
(0,2) (0,2) yes
Table 3: The incompatibility status for all possible coarse-graining (CG) of outcomes for Alice’s measurements ($A_{1}$ and $A_{2}$) as inferred from the statistics of CGLMP experiment. Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurement remains incompatible. On the other hand, “????” denotes that we can not make any conclusion about their incompatibility.
IV.2 Convex mixing of measurements
Device-independent witness
Here, one can simply use the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al. (2009). Considering the convex mixture of measurements for any three binary-outcome measurements, this result can be used to witness full incompatibility w.r.t. convex mixing of measurements.
Semi-device-independent witness
For three or more numbers of measurements we can witness whether the measurements are fully-incompatible w.r.t convex-mixture in a semi-device independent manner by constructing a suitable random access codes task.
Theorem 4.
Three noisy Pauli measurements of Eq.(16) are witnessed to be fully incompatible w.r.t convex-mixing via RAC iff: $\sqrt{\frac{2}{3}}<\nu\leq 1.$
Proof.
Let us denote the measurements of Eq.(16) as $\mathcal{M}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{0}\sigma_{x}\right)$, $\mathcal{M}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{0}\sigma_{x}\right)$, $\mathcal{N}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{1}\sigma_{y}\right)$, $\mathcal{N}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{1}\sigma_{y}\right)$, $\mathcal{R}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{2}\sigma_{z}\right)$, and $\mathcal{R}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{2}\sigma_{z}\right)$, and consider the measurements $Q^{i}$ with $i\in\{1,2\}$, to be formed by the convex-mixing of $\mathcal{M}$ and $\mathcal{N}$ as
$Q^{1}=\{p\,\,M_{0}+(1-p)\,\,N_{0},\,\,pM_{1}+(1-p)N_{1}\},$
$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$
In order to witness the incompatibility status of $(Q^{i},R)$, one can construct a (2,2,2) RAC task. Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $Q^{i}$ which can be realized by the convex mixing of measurements $\mathcal{M}$ and $\mathcal{N}$ as follows: Bob tosses a biased coin which gives a ‘Head’ with probability $p$ and a ‘Tail’ with probability $(1-p)$. On obtaining ‘Head’, Bob performs measurement $\mathcal{M}$, and for ‘Tail’ he performs measurement $\mathcal{N}$. For the second bit, he performs measurement $\mathcal{R}$. The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task (considering $\nu_{0}=\nu_{1}=\nu_{2}$) is:
$P(2,2,2)$ $=$ $\frac{1}{8}\sum_{j,k=0}^{1}||Q^{i}_{j}+R_{k}||$ (62)
$=$ $\frac{1}{4}(2+\sqrt{2\nu^{2}-2p\nu^{2}+2p^{2}\nu^{2}}).$
It can be shown that for all possible convex-mixing the maximum average success probability is the same as Eq.(62). Now, to be fully incompatible w.r.t convex-mixing $P(2,2,2)>\frac{3}{4},\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(62)). Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{\frac{2}{3}}$ the measurements are fully incompatible w.r.t convex-mixing.
In Observation 4, we have found that this kind of three qubit measurements with noise become fully incompatible w.r.t convex mixture for $\nu>\sqrt{\frac{2}{3}}$. So, we can conclude that when these three measurements are fully incompatible w.r.t convex mixing, quantum advantage is obtained in a suitable Random Access Code task.
[…]
Examples.— Consider the following three rank-one projective measurements acting on $\mathbb{C}^{3}$:
$X=\{|0_{x}\rangle\langle 0_{x}|,|1_{x}\rangle\langle 1_{x}|,|2_{x}\rangle\langle 2_{x}|\},$
$Y=\{|0_{y}\rangle\langle 0_{y}|,|1_{y}\rangle\langle 1_{y}|,|2_{y}\rangle\langle 2_{y}|\},$
and $Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (63)
where,
$|0_{x}\rangle$ $=$ $\frac{1}{2}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle+\frac{1}{2}|2\rangle,$
$|1_{x}\rangle$ $=$ $-\frac{1}{\sqrt{2}}|0\rangle+\frac{1}{\sqrt{2}}|2\rangle,$
$|2_{x}\rangle$ $=$ $\frac{1}{2}|0\rangle-\frac{1}{\sqrt{2}}|1\rangle+\frac{1}{2}|2\rangle,$
$|0_{y}\rangle$ $=$ $-\frac{\mathbbm{i}}{2}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle+\frac{\mathbbm{i}}{2}|2\rangle,$
$|1_{y}\rangle$ $=$ $\frac{\mathbbm{i}}{\sqrt{2}}|0\rangle+\frac{\mathbbm{i}}{\sqrt{2}}|2\rangle,$
$|2_{y}\rangle$ $=$ $-\frac{\mathbbm{i}}{2}|0\rangle-\frac{1}{\sqrt{2}}|1\rangle+\frac{\mathbbm{i}}{2}|2\rangle,$
$|0_{z}\rangle$$\equiv$ $|0\rangle,|1_{z}\rangle\equiv|1\rangle,|2_{z}\rangle\equiv|2\rangle.$ (64)
We take the following convex mixture of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$. We now consider the $(2,3,3)$ RAC game involving the two measurements $A$ and $Z$. The maximum average success probability $P(2,3,3)$ for this RAC task is given by Carmeli et al. (2020),
$P(2,3,3)=\frac{1}{18}\sum_{i,j=0}^{2}||A_{i}+Z_{j}||,$ (65)
which turns out to be greater than $\frac{2}{3}$ i.e. $P_{CB}(2,3,3)$ for all $p\in[0,1]$. Hence, the measurements $A$ and $Z$ are incompatible.
Next, consider the following measurement (taking an arbitrary convex mixture of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$. Note that the aforementioned measurement $A$ is nothing but $A_{0,0,1,1,2,2}$ in the present notation. Following a similar calculation, it can be shown that any such $A_{i,j,k,l,m,n}$ is incompatible with $Z$ for all $p\in[0,1]$.
Similarly, taking an arbitrary convex combination of $X$ and $Z$ (or, $Y$ and $Z$), it can be shown that the new measurement is incompatible with $Y$ (or, $X$) for all $p\in[0,1]$.
A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements: $M=\{M_{0},M_{1},\ldots,M_{n-1}\}.$ We make $m$ partitions $S=\{S_{i}\}_{i=0}^{m-1}$, $\bigcup_{i}S_{i}=M,\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $m$ measurements produced from the convex mixing of $n$ measurements from each of the $m$ partitions. We can construct a $(m,d,d)$ RAC game where Alice has an $m$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{m-1})$. Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in\{0,1,\ldots,m-1\}$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$. He performs the measurement which is obtained by the convex mixtures of the measurements from the partition $S_{y}$ and declares the outcome of the measurement. Now, if the success probability $P(m,d,d)$ is greater than $P_{CB}(m,d,d)$ Saha et al. (2023) for all possible convex mixtures, we can operationally detect that the $m$ partitioned measurements are incompatible under convex mixing.
V Conclusions
Measurement incompatibility is the spectacle in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems. It is one of the fundamental ingredients for non-classical correlations and merits of quantum information science. Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases. Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance both from the foundational and practical perspectives. In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and convex-mixing of different measurements.
Through our present analysis we have investigated the different levels of incompatibility arising under the above classical operations. Since environmental effects are ubiquitous in practical scenarios, here we have investigated the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations. Furthermore, we have developed a method to operationally witness different levels of measurement incompatibility in the device-independent framework involving Bell-type experiments, and also in the semi-device-independent framework involving prepare-and-measure experiments.
Several examples have been provided to illustrate the efficacy of the operational witnesses proposed here. Our results are particularly useful for the purpose of comparing measurements in terms of their degree of incompatibility. For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required.
Our present study motivates future work on several open directions. The condition for full incompatibility for two projective measurements with respect to coarse-graining of outcomes can be extended for more number of projective measurements. Similarly, the criterion for full incompatibility of projective measurements with respect to convex mixing could be generalized for any set of measurements. Further investigation would be required to see if device-independent witnesses could be formulated for more than three outcome measurements with respect to coarse-graining and more than three measurements with repect to convex mixing. It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from a single experimental statistics.
VI Acknowledgements
[…]